OpenStudy (anonymous):

Which number makes this equation true? v^2 + 10v + 16 = (v + 8)(v + ?)

OpenStudy (anonymous):

@jim_thompson5910

Parth (parthkohli):

I know a shortcut... can I say it?

jimthompson5910 (jim_thompson5910):

if you were to set up a table, the lower right corner would have the value of 8*? where the answer would replace the question mark

OpenStudy (anonymous):

yes please @ParthKohli

jimthompson5910 (jim_thompson5910):

8*? has to equal 16 so basically 8p = 16 p = ???

jimthompson5910 (jim_thompson5910):

where p is that answer

OpenStudy (anonymous):

2

jimthompson5910 (jim_thompson5910):

yep, you nailed it

OpenStudy (anonymous):

yayyy :)

jimthompson5910 (jim_thompson5910):

so v^2 + 10v + 16 = (v + 8)(v + ?) turns into v^2 + 10v + 16 = (v + 8)(v + 2)

OpenStudy (anonymous):

What is the factored form of x^2 + 6x + 8?

Parth (parthkohli):

The shortcut says that in \(v^2 + \color{blue}{10}v + \rm something \) and \((v + \color{green}{a})(v + \color{green}{b})\), the sum of the greens is the blue guy.

jimthompson5910 (jim_thompson5910):

find two numbers that a) multiply to 8 (last term) AND b) add to 6 (middle coefficient)

OpenStudy (anonymous):

im confused @ParthKohli

OpenStudy (anonymous):

2 and 4 . @jim_thompson5910

Parth (parthkohli):

It's simple. You have \((v + 8)(v + ?)\). This shortcut says that \(8 + ? = 10\).

jimthompson5910 (jim_thompson5910):

So x^2 + 6x + 8 factors to (x+2)(x+4) The rule is that x^2 + bx + c factors to (x+m)(x+n) where m+n = b and m*n = c

OpenStudy (anonymous):

ohhh! @ParthKohli thanks :)

jimthompson5910 (jim_thompson5910):

So it's just a matter of finding those two numbers then writing it in the form (x+number1)(x+number2)

Parth (parthkohli):

I know that was sarcastic. -_-

OpenStudy (anonymous):

hold on im thinking @jim_thompson5910

OpenStudy (anonymous):

5 and 3?? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

you found the numbers already by saying 2 and 4 2+4 = 6 (middle coefficient) 2*4 = 8 (last term)

OpenStudy (anonymous):

ohh.

jimthompson5910 (jim_thompson5910):

so because you found two numbers that multiply to 8 AND add to 6 the factorization of x^2 + 6x + 8 would be (x+2)(x+4)

OpenStudy (anonymous):

so it's (x + 4)(x + 2) ?

jimthompson5910 (jim_thompson5910):

go ahead and expand out (x+2)(x+4) you'll find that you'll get x^2 + 6x + 8 back again

jimthompson5910 (jim_thompson5910):

so essentially what you've done is gone in reverse of expanding (which was done in the previous problem)

OpenStudy (anonymous):

i mostly get what you're saying. i have one more if you're willing to help?

jimthompson5910 (jim_thompson5910):

correct, the order of the factors doesn't matter so (x+2)(x+4) is the same as (x+2)(x+4)

jimthompson5910 (jim_thompson5910):

i meant (x+2)(x+4) is the same as (x+4)(x+2)

jimthompson5910 (jim_thompson5910):

sure go for it

OpenStudy (anonymous):

What is the factored form of x^2 – 7x + 12? and dont worry i knew what you meant lol

jimthompson5910 (jim_thompson5910):

same idea, just different numbers

jimthompson5910 (jim_thompson5910):

find two numbers that multiply to 12 AND add to -7

OpenStudy (anonymous):

-3 and -4

jimthompson5910 (jim_thompson5910):

perfect

jimthompson5910 (jim_thompson5910):

so it factors to (x-3)(x-4)

OpenStudy (anonymous):

(x – 4)(x – 3) ?

jimthompson5910 (jim_thompson5910):

unfortunately this trick only works if the leading coefficient is 1, but it's still useful

jimthompson5910 (jim_thompson5910):

you got it

jimthompson5910 (jim_thompson5910):

that's the same as (x-3)(x-4) because the order of the factors does not matter ex: 2 time 3 = 6 and 3 times 2 = 6

OpenStudy (anonymous):

okay i have a different kind of problem... The area of a garden is given by the trinomial g2 – 2g – 24. The garden’s length is g + 4. What is the garden’s width?

jimthompson5910 (jim_thompson5910):

well it's actually the same type of problem, just worded differently

jimthompson5910 (jim_thompson5910):

they want you to factor g^2 - 2g - 24

jimthompson5910 (jim_thompson5910):

what do you get when you do?

OpenStudy (anonymous):

write me an equation then and let me solve it :)

OpenStudy (anonymous):

like for the question

jimthompson5910 (jim_thompson5910):

find me two numbers that multiply to -24 and add to -2 at the same time

OpenStudy (anonymous):

-6 and -4

jimthompson5910 (jim_thompson5910):

-6 plus -4 = -10 and it does NOT equal -2

jimthompson5910 (jim_thompson5910):

so try again

OpenStudy (anonymous):

oops lol hold on

OpenStudy (anonymous):

-6 and 4? -6*4= 24 -6+4 = -2 i think lol

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

So g^2 - 2g - 24 factors to (g-6)(g+4)

jimthompson5910 (jim_thompson5910):

notice how g+4 is a factor

jimthompson5910 (jim_thompson5910):

the other factor is g-6

jimthompson5910 (jim_thompson5910):

let's go back to the question The area of a garden is given by the trinomial g2 – 2g – 24. The garden’s length is g + 4. What is the garden’s width?

OpenStudy (anonymous):

g-6!

jimthompson5910 (jim_thompson5910):

correct since Area = Length times Width

OpenStudy (anonymous):

:D

jimthompson5910 (jim_thompson5910):

so if you have some area expression and you factor it, you get (Length)*(Width) where length and width are the factors of the area

OpenStudy (anonymous):

I actually get that!

jimthompson5910 (jim_thompson5910):

it only works for rectangles though

jimthompson5910 (jim_thompson5910):

that's great

OpenStudy (anonymous):

hb this one? What is the factored form of 4x^2 + 12x + 5?

jimthompson5910 (jim_thompson5910):

ok now the leading coefficient isn't 1 anymore, so we can't use that trick (well not the whole thing anyways)

jimthompson5910 (jim_thompson5910):

but we can still use some of the trick

jimthompson5910 (jim_thompson5910):

first multiply 4 (first coefficient) and 5 (last term) to get 4*5 = 20

jimthompson5910 (jim_thompson5910):

now list two numbers that multiply to 20 AND add to 12 (middle coefficient)

OpenStudy (anonymous):

2 and 10

OpenStudy (anonymous):

is it (4x + 5)(x + 1) ?

jimthompson5910 (jim_thompson5910):

so we break 12x into 2x + 10x then we factor by grouping 4x^2+12x+5 4x^2+2x+10x+5 (4x^2+2x)+(10x+5) 2x(2x+1)+(10x+5) 2x(2x+1)+5(2x+1) (2x+5)(2x+1)

jimthompson5910 (jim_thompson5910):

hopefully factoring by grouping is a familiar term

OpenStudy (anonymous):

so the final answer is (2x + 1)(2x + 5), right?

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

hopefully you see how I got all that

OpenStudy (anonymous):

okay, thanks :D and lol yeah mostly.

OpenStudy (anonymous):

The area of a rectangular swimming pool is 10x2 – 19x – 15. The length of the pool is 5x + 3. What is the width of the pool?

jimthompson5910 (jim_thompson5910):

alright, just keep practicing and it'll make more sense

jimthompson5910 (jim_thompson5910):

10*(-15) = -150

jimthompson5910 (jim_thompson5910):

find two numbers that multiply to -150 and they add to -19

OpenStudy (anonymous):

why do you multiply 10 and -15?

jimthompson5910 (jim_thompson5910):

because I'm using the AC method for factoring this method starts out by multiplying the first and last coefficients

jimthompson5910 (jim_thompson5910):

then you find two numbers that multiply to that result and add to the middle number

OpenStudy (anonymous):

so the formula is ax^2 + bx + c right?

jimthompson5910 (jim_thompson5910):

that's a basic form of a quadratic, yes

jimthompson5910 (jim_thompson5910):

you would multiply 'a' and 'c' to get a*c or just ac then you find two numbers that multiply to ac and add to b

jimthompson5910 (jim_thompson5910):

so that's why they call it the ac method

OpenStudy (anonymous):

I cant find the two numbers that multiply to -150 and add to -19..

OpenStudy (anonymous):

im too dumb..

jimthompson5910 (jim_thompson5910):

no you're not

jimthompson5910 (jim_thompson5910):

stop that

jimthompson5910 (jim_thompson5910):

there's a long list to check, but luckily there's a shortcut

jimthompson5910 (jim_thompson5910):

have you heard of the quadratic formula before?

OpenStudy (anonymous):

noo. my teacher sucks this year. she is retiring so she doesnt teach us anything useful, she just assigns stuff and tells us to read our book to figure out how to do it, but the book has no tips or tricks.